Starting from rest, a car accelerates at a constant rate of 2.6 m/s2 for a time of 5 seconds.
(a) Compute the velocity at 1 second?
(b) Compute the velocity at 2 seconds?
(c) Compute the velocity at 3 seconds?
(d) Compute the velocity at 4 seconds?
(e) Compute the velocity at 5 seconds?
(f) Compute the distance traveled at 1 second?
(g) Compute the distance traveled at 2 seconds?
(h) Compute the distance traveled at 3 seconds?
(i) Compute the distance traveled at 4 seconds?
(j) Compute the distance traveled at 5 seconds
Answer #1

vi = 0 and a = 2.6 m/s^2

and v = vi + a t and d = vi t + a(t^2)/2


(A) v = 0 + (2.6)(1) = 2.6 m/s


(B) v = 0 + (2.6)(2) = 5.2 m/s


(C) v = 7.8 m/s

(D) v = 10.4 m/s

(E) 13 m/s


(f) d = 0 + (2.6)(1^2)/2 = 1.3 m


(g) d = 0 + (2.6)(2^2)/2 = 5.2 m


(h) d = 0 + (2.6)(3^2)/2 = 11.7 m

(i) d = 20.8 m

(j) d = 32.5 m